Write the slope intercept form of the equation of the line through the given point, with the given slope.
(0), (-5) slope + 0

Answer by hailee.devoll(2)

y=mx+b. -5=0(0)+b so -5=b then plug it back in and you get y=-5 because 0 times something is 0. So your answer is y=-5